# Remixing Music for Hearing Aids Using Ensemble of Fine-Tuned Source Separators

### 11 Jan 2024 v1

@@ Abstract
Index Terms—  Cadenza ICASSP 2024 Grand Challenge, hearing aids, music enhancement, music source separation, remixing
@@ 1 Introduction
The Cadenza challenge provides a data set for training, a validation data set, and a test data set for the final evaluation. The data includes songs, sets of gains for the VDBO components, and listener audiograms. An important aspect of the data is that the audio signals are convolved with head-related transfer functions (HRTFs) to simulate a person listening to stereo speakers in a free field with different head positions. This introduces cross-talk to the stereo channels, meaning each received stereo channel contains a combination of the separate transmitted stereo channels, depending on the head position relative to the speakers.
@@
Our proposed ensemble system placed first in the Cadenza ICASSP 2024 Grand Challenge, with the best average HAAQI score on the evaluation data set. This paper describes the system and presents evaluation results demonstrating its effectiveness in the challenge. We analyze the results and use ablation studies the measure the importance of different system components.
@@ 2 System Design
For the DTTNet and KUIELab-MDX-Net models, which required a separate model trained for each VDBO track, we selected salient segments from the training set using the source activity detection algorithm mentioned in [8].

### 01 Feb 2024 v2

@@ Abstract
© IEEE 2024. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. Index Terms—  Cadenza ICASSP 2024 Grand Challenge, hearing aids, music enhancement, music source separation, remixing
@@ 1 Introduction
The Cadenza challenge provides a data set for training, a validation data set, and a test data set for the final evaluation. The data includes songs, sets of gains for the VDBO components, and listener audiograms. An important aspect of the data is that the audio signals are convolved with head-related transfer functions (HRTFs) to simulate a person listening to stereo speakers in a free field with different head positions. This introduces crosstalk to the stereo channels, meaning each received stereo channel contains a combination of the separate transmitted stereo channels, depending on the head position relative to the speakers.
@@
Our proposed ensemble system placed first in the Cadenza ICASSP 2024 Grand Challenge, with the best average HAAQI score on the evaluation data set. This paper describes the system and presents evaluation results demonstrating its effectiveness in the challenge. We analyze the results and use ablation studies to measure the importance of different system components.
@@ 2 System Design
The KUIELab-MDX-Net and DTTNet architectures are both based on a Time-Frequency Convolutions with Time-Distributed Fully-connected networks UNet (TFC-TDF UNet) architecture [9]. Both of these architectures require a separate model trained for each VDBO track. Each KUIELab-MDX-Net model contains 7.4 million parameters, and each DTTNet model contains 5 million parameters. The HDemucs model uses a convolutional encoder-decoder architecture that operates on both frequency-domain and time-domain representations. The HDemucs architecture requires a single model to separate the four VDBO tracks, and it contains 83.6 million parameters.
@@
For the DTTNet and KUIELab-MDX-Net models we selected salient segments from the training set using the source activity detection algorithm mentioned in [8].